What are the prices of a call option and a put option with the following characteristics? (Do not round intermediate calculatPlease Show all work and formulas
Answer #1

stock price exercise price risk free interest rate time to expiration standard deviation (volatality) 0.33 0.53 85 4.0% CÓ 5put price= Xen-rt N(-d2) - S.N(-d1) d1 = d2 = di-v* T^42 In(so/X) + (r+(v^2/2))T v* T^V2 d1 = 0.10613511 0.305995643 0.346851